1. how many square units are in the area of the circle described by x^2+y^2=25? expres your answer in terms of pi.
Answer by Nate(3500)

(x - h)^2 + (y - k)^2 = r^2
where (h,k) is the vertex, and (r) is the radius
r^2 = 25
r = 5
This is a circle with a radius of five units.
Area = (pi)r^2
Area = 25(pi)